OpenStudy (anonymous):

If 1,000 students take a quiz that has a mean of 40 minutes, a standard deviation of 8 minutes, and is normally distributed, how many would you expect to finish in between 24 and 48 minutes?

OpenStudy (anonymous):

\[\begin{align*} P(24<X<48)&=P\left(\frac{24-40}{8}<\frac{X-40}{8}<\frac{48-40}{8}\right)\\ &=P(-2<Z<1)\\ &\approx0.816 \end{align*}\] according to WA: http://www.wolframalpha.com/input/?i=P%28-2%3CZ%3C1%29 So what you have here is the proportion of students that fall in the desired interval. To find the actual number of students, multiply this proportion by the total number of students: \[1000\times0.816=816\text{ students}\]